OpenStudy (anonymous):

The table below shows the distance y, in miles, traveled by a toy car in x minutes: Time (x) (minutes) 10 20 30 40 Distance (y) (miles) 4 8 12 15 Part A: What is the most likely value of the correlation coefficient of the data in the table? Based on the correlation coefficient, describe the relationship between time and distance traveled by the toy car. (4 points) [Choose the value of correlation coefficient from 1, 0.99, 0.5, 0.02] Part B: What is the value of the slope of the graph of distance versus time between 10 minutes and 30 minutes and what does the slope represent? (3 poin

OpenStudy (anonymous):

@kagıtucak Help please?

OpenStudy (kagıtucak):

i dont know part A

OpenStudy (kagıtucak):

but part b slope=differences between y / differences between x slope=8/20=2/5

OpenStudy (anonymous):

Okay thank you there is apart C Part C: Does the data in the table represent correlation or causation? Explain your answer. @kagıtucak

OpenStudy (kagıtucak):

i dont know what correlation and correlation coefficient is sorry

OpenStudy (anonymous):

Okay thank you I'll figure it out, could you help me with another question? @kagıtucak

OpenStudy (kagıtucak):

write it

OpenStudy (anonymous):

Stacey is studying the sale of a particular brand of cereals from the year 1995 to 2004. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t) = t2 + 6t + 48 Part A: What does the y-intercept of the graph of the function represent? (4 points) Part B: What is the reasonable domain of the graph of the function? (3 points) Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work. @kagıtucak

OpenStudy (kagıtucak):

for y intercept take t=0 so y=48

OpenStudy (kagıtucak):

Part B: there is no value which make the s(t) 0 becuase its delta is 36-4*1*48<0 so everywhere is its domain

OpenStudy (kagıtucak):

Part C: for average rate take derivative of the function it is 2t+6 t=1 it is 8 t=4 it is 14 14-8=6